OpenStudy (anonymous):

Media services charges $30 for for a phone and $20/month for its economy plan. The equation c=20t+30 describes the total cost, c, of operating a media Services phone for t months... the total cost for 3 months service is ??????(pretty sure this is $90) If the costumer has only $80 how many months of service can she recieve a) 2 1/2 B) 3 c) 2 1/3 d) 2

OpenStudy (anonymous):

a) 2 1/2............solve 80=20t+30....................t=50/20=2.5

OpenStudy (saifoo.khan):

still stuck?

OpenStudy (anonymous):

uhm... no I think i got it.. 2.5=2 and 1/2 so yeah pretty much self explanatory lol

OpenStudy (saifoo.khan):

No, with the further graphing part.

OpenStudy (anonymous):

oh yeah no i dont know hwo to graph it h ah a

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

wait. which week is that?

OpenStudy (anonymous):

week 9 finals HAHAHA

OpenStudy (saifoo.khan):

or if you can post a photo then i can tell u the answer.

OpenStudy (saifoo.khan):

lol, best of luck.. 66 questions will take a while.

OpenStudy (anonymous):

yeah im halfway done been working on it since about 8 this morning

OpenStudy (saifoo.khan):

lol, what do u think u can get? a A? B?

OpenStudy (saifoo.khan):

my wishes are with you... I hope you achieve an A. :D

OpenStudy (anonymous):

thank you :D

OpenStudy (saifoo.khan):

no probs! :D

OpenStudy (anonymous):

I will probably fail miserably... I did the last time and had to re take the course blah

OpenStudy (saifoo.khan):

haha, if you want help, i can help you through. :)